# Residual Diffusion Bridge Model for Image Restoration

### v1

@@ 5.3 Comparative Experiments
Figure 6: The visualization results of zero-shot generalization in real-world TOLED dataset. Zoom in for best view.
@@ 5.4 Ablation Study
To thoroughly explore the efficacy of our method, we carry out ablation studies encompassing three distinct categories: Influence of various implementation configurations. Our RDBM formulations are governed by the schedule $\{\theta_{t}\}$ and stationary variance $\lambda$. Initially, we adopt the empirical choice $\lambda\!=\!\frac{10}{255}$ [72] and compare performance across different noise schedules, as reported in Tab. 3. It is evident that the optimal noise schedules differ for distinct restoration tasks, with the cosine schedule generally yielding the best results. Building on this finding, we further conduct a quantitative comparison among diverse stationary variance $\lambda$, as presented in Tab. 4. The results indicate that $\lambda\!=\!\frac{10}{255}$ with cosine noise schedule is the optimal configuration.
@@
Restomer [75] 11.500 0.445 0.077 0.494 11.094 0.495 0.106 0.330
AirNet [28] 5.705 0.103 0.324 1.072 9.706 0.430 0.117 0.403
Prompt-IR [50] 11.589 0.429 0.075 0.541 13.088 0.504 0.105 0.336
ProRes [43] 10.284 0.433 0.103 0.473 28.452 0.834 0.002 0.212
IDR [79] 13.583 0.466 0.057 0.551 24.259 0.759 0.008 0.253
IRSDE [40] 16.983 0.615 0.029 0.475 27.163 0.811 0.002 0.243
AutoDIR [20] 8.627 0.404 0.151 0.406 9.354 0.443 0.130 0.338
DA-CLIP [41] 16.788 0.559 0.025 0.469 27.256 0.789 0.003 0.201
GOUB [72] 12.922 0.525 0.053 0.446 23.177 0.761 0.007 0.269
ConvIR [11] 9.370 0.429 0.130 0.477 13.659 0.558 0.091 0.316
DeepSNNet [12] 10.195 0.411 0.113 0.534 17.394 0.576 0.075 0.303
AWRaCLe [52] 11.208 0.431 0.091 0.513 10.540 0.495 0.121 0.331
MaIR [30] 11.072 0.423 0.086 0.529 23.637 0.770 0.005 0.271
@@
Figure 7: The visualization results of zero-shot generalization in real-world TOLED dataset. Zoom in for best view.
@@ 5.5 Zero-Short Real-world Generation
Unknown task generalization. POLED and TOLED [83] are captured by under-display cameras in high-resolution with different degradation types, which fully meet the real-world scene. The quantitative results are reported in Tab. 7 while the visual comparisons are illustrated in Fig. 6. Evidently, our method achieves the best metric evaluation and our restored image is the most similar to ground-truth.
Known task generalization. As real-world datasets mainly have no ground truth, we use the non-reference metric, i.e., MetaIQA [84] and NIQE [47], to assess the perceptual quality, as provided in Tab. 8. Results show that our method outperforms other universal models in various benchmarks.
@@
Restomer [75] 0.231 13.115 0.328 3.828 0.357 5.845 0.437 4.400 0.303 6.734
AirNet [28] 0.232 11.668 0.280 3.674 0.347 6.091 0.440 4.623 0.286 6.393
Prompt-IR [50] 0.232 11.439 0.308 3.797 0.361 5.840 0.437 4.962 0.286 6.670
ProRes [43] 0.226 13.110 0.348 3.933 0.355 5.976 0.434 5.444 0.297 6.574
IDR [79] 0.231 11.100 0.324 3.866 0.363 5.850 0.453 4.634 0.300 6.683
IRSDE [40] 0.230 11.391 0.351 3.809 0.357 5.874 0.427 4.134 0.285 6.289
AutoDIR [20] 0.231 10.800 0.366 3.910 0.373 5.831 0.470 9.881 0.308 6.493
DA-CLIP [41] 0.232 10.604 0.334 3.720 0.361 5.864 0.460 6.531 0.310 6.058
GOUB [72] 0.231 11.566 0.373 3.928 0.360 5.853 0.458 4.104 0.278 6.303
ConvIR [11] 0.236 10.280 0.370 3.723 0.364 5.813 0.446 4.645 0.313 6.465
DeepSNNet [12] 0.231 11.446 0.348 3.896 0.367 5.882 0.436 4.662 0.301 6.525
AWRaCLe [52] 0.232 12.016 0.366 3.796 0.363 5.898 0.426 4.649 0.306 6.516
MaIR [30] 0.234 10.804 0.350 3.666 0.363 5.890 0.245 22.446 0.284 6.590
@@ 5.7 Image Translation and Inpainting
RDBM owns distinct advantages in mapping the data distribution to the prior distribution, thereby enabling extensive validation on similar computer vision tasks. To this end, we expand our experimental settings on image-to-image translation and image inpainting to fully demonstrate the potential of our method. The former aims to transform an input image from one domain to another while preserving certain essential semantic or structural features. The latter focuses on filling in missing or corrupted regions within an image. Specifically, we adopt the widely used edge2handlebags [17] dataset for image-to-image translation and apply the Celebrate-HQ dataset [21] with masks provided in  [33] for image inpainting. All these datasets are scaled to 256 × 256. We additionally employ Fréchet Inception Distance (FID) [14] for evaluation. Qualitative comparisons and quantitative results are presented in Fig. 7 and Tab. 9, respectively. Clearly, our method achieves the best visual effects and the best metrics assessments.
@@
Method Image Translation [17] Image Inpainting [21]
@@
DDPM [61] 8.39 0.447 8.39 0.412 19.22 0.746 0.526 0.126
ReFlow [37] 10.46 0.442 5.76 0.374 23.53 0.822 0.307 0.149
BBDM [29] 14.75 0.635 7.46 0.248 20.36 0.653 0.386 0.169
I2SB [34] 12.57 0.615 6.64 0.357 27.34 0.890 0.379 0.054
RDDM [36] 14.66 0.645 5.72 0.256 23.94 0.852 0.167 0.119
GOUB [72] 16.58 0.700 8.76 0.288 31.56 0.920 0.321 0.065
@@ Appendix A Doob’s $h$ transform
Proof: In theory, $p(\mathbf{x}_{t}\mid\mathbf{x}_{0})$ and $p(\mathbf{x}_{T}\mid\mathbf{x}_{t})$ satisfy the Kolmogorov Forward Equation (KFE) and Kolmogorov Backward Equation (KBE), respectively [54], as formulated below:
@@ Appendix E Process of Reverse Inference
$\mathbf{x}_{t-1}=\kappa_{t}\mathbf{x}_{t}+\eta_{t}\boldsymbol{\mu}+\gamma_{t}\mathbf{x}_{0}+\sigma_{t}\boldsymbol{\pi}\epsilon_{t}$ (E.13)
$=\kappa_{t}(\boldsymbol{\mu}+(\mathbf{x}_{0}-\boldsymbol{\mu})\Theta_{t}+\boldsymbol{\pi}\Sigma_{t}\epsilon_{t})+\eta_{t}\boldsymbol{\mu}+\gamma_{t}\mathbf{x}_{0}+\sigma_{t}\boldsymbol{\pi}\epsilon_{t}$
$=(\kappa_{t}+\eta_{t}-\kappa_{t}\Theta_{t})\boldsymbol{\mu}+(\kappa_{t}\Theta_{t}+\gamma_{t})\mathbf{x}_{0}+\boldsymbol{\pi}(\kappa_{t}^{2}\Sigma_{t}^{2}+\sigma_{t}^{2})^{\frac{1}{2}}\epsilon_{t},$
@@
$\Sigma_{t-1}^{2}$ $=\kappa_{t}^{2}\Sigma_{t}^{2}+\sigma_{t}^{2}.$ (E.16)
By setting $\sigma_{t}=0$:
@@ Appendix F Connections Among Existing Diffusion Bridge
$d\mathbf{x}_{t}=\theta_{t}\coth(\overline{\theta}_{t:T})(\boldsymbol{\mu}-\mathbf{x}_{t})dt+\sqrt{2\pi^{2}\lambda\theta_{t}}d\omega_{t},$ (F.2)
@@ F.1 Connections to Variance-Exploding (VE) and Variance-Preserving (VP) SDEs
SMLD [61] primarily introduces two mainstream diffusion formulations, namely VP and VE. For a given generalized OU process in Eq. (F.1), there exists relationships:
@@
where we set the $\theta_{t}\rightarrow\sigma_{t}^{2}$, implying $\lambda=1$. On this basis, DDBM [82] further extends such diffusion configuration to bridge models, which are also special cases of our formulation under specific configurations:
@@ F.2 Connections to Brownian Bridge SDEs
Brownian bridge is a fundamental architecture for diffusion model, which are widely adopted in BBDM [29], I2SB [34]. By setting $\theta_{t}\rightarrow 0$ with condition $2\lambda\theta_{t}=1=\sigma_{t}^{2}$, we can derive the Brownian Bridge as formulated below:
@@
$=\lim\limits_{\theta_{t}\rightarrow 0,\sigma_{t}^{2}\rightarrow 1}^{\boldsymbol{\pi}=1}\{d\mathbf{x}_{t}=\frac{\boldsymbol{\mu}-\mathbf{x}_{t}}{T-t}dt+\sigma_{t}d\omega_{t}\}$
@@ F.3 Connections to Flow Matching
Flow-based generative models [32, 37] design a deterministic probability path that linearly interpolates between a prior and the data distribution, and then directly learn a time-dependent vector field whose integral trajectories realize this path. By discarding the stochastic noise ($\sigma_{t}=0$) and adopting the Brownian bridge configuration, Eq. (F.2) can be transformed into:
@@ F.4 Connections to OU Bridge SDEs
Eq. (F.2) can be transformed into naive OU bridge [72] by setting $\pi=1$ to recover global noise perturbation:
@@ F.5 Connections to Stochastic Interpolants
Stochastic interpolants [1] define a unified framework for flows and diffusions, which can be expressed as:
@@ H.1 Summary about the Datasets
Table S1: Summary of the natural image restoration datasets utilized in this paper.
@@
Deraining DID [77] Synthetic - 1,200
Rain13K [19] Synthetic 13,711 -
Rain_100 [71] Synthetic - 200
DeRaindrop [51] Real 861 307
GT-Rain [4] Real 26,125 2,100
RealRain-1k [31] Real 1792 448
Low-light Enhancement LOL [67] Real 485 15
@@
VE-LOL-L [35] Synthetic/Real 900/400 100/100
NPE [65] Real - 8
DICM [26] Real - 64
Desnowing CSD [6] Synthetic 8,000 2,000
Snow100K-Real [38] Real - 1,329
Dehazing SOTS [27] Synthetic - 500
ITS_v2 [27] Synthetic 13,990 -
D-HAZY [10] Synthetic 1,178 294
NH-HAZE [3] Real - 55
Dense-Haze [2] Real - 55
NHRW [78] Real - 150
Deblur GoPro [48] Synthetic 2,103 1,111
RealBlur [53] Real 3,758 980
@@ H.3 More Visual Comparisons on Real-world Scene Generalization
Unknown task generalization. Unknown task image restoration is performed on both POLED and TOLED [83]. Visual comparisons on the POLED dataset are provided in Fig. S10. The results show that our method can generalize to real-world scenes and achieve competitive visual results.
Figure S10: The visualization results of zero-shot generalization in real-world POLED dataset. Zoom in for best view.
@@ H.4 More Visual Comparisons on Image Translation and Inpainting
Figure S11: The visualization results of image translation. Zoom in for best view.
Figure S12: The visualization results of image inpainting. Zoom in for best view.

### v2

@@ 5.3 Comparative Experiments
Figure 6: Visualization results of zero-shot generalization in real-world TOLED dataset. Zoom in for best view.
@@ 5.4 Ablation Study
To thoroughly explore the efficacy of our method, we carry out ablation studies encompassing three distinct categories: Influence of various implementation configurations. Our RDBM formulations are governed by the schedule $\{\theta_{t}\}$ and stationary variance $\lambda$. Initially, we adopt the empirical choice $\lambda\!=\!\frac{10}{255}$ [yue2024image] and compare performance across different noise schedules, as reported in Tab. 3. It is evident that the optimal noise schedules differ for distinct restoration tasks, with the cosine schedule generally yielding the best results. Building on this finding, we further conduct a quantitative comparison among diverse stationary variance $\lambda$, as presented in Tab. 4. The results indicate that $\lambda\!=\!\frac{10}{255}$ with cosine noise schedule is the optimal configuration.
@@
Restomer [zamir2022restormer] 11.500 0.445 0.077 0.494 11.094 0.495 0.106 0.330
AirNet [li2022all] 5.705 0.103 0.324 1.072 9.706 0.430 0.117 0.403
Prompt-IR [potlapalli2023promptir] 11.589 0.429 0.075 0.541 13.088 0.504 0.105 0.336
ProRes [ma2023prores] 10.284 0.433 0.103 0.473 28.452 0.834 0.002 0.212
IDR [zhang2023ingredient] 13.583 0.466 0.057 0.551 24.259 0.759 0.008 0.253
IRSDE [luo2023image] 16.983 0.615 0.029 0.475 27.163 0.811 0.002 0.243
AutoDIR [jiang2024autodir] 8.627 0.404 0.151 0.406 9.354 0.443 0.130 0.338
DA-CLIP [luo2024controlling] 16.788 0.559 0.025 0.469 27.256 0.789 0.003 0.201
GOUB [yue2024image] 12.922 0.525 0.053 0.446 23.177 0.761 0.007 0.269
ConvIR [cui2024revitalizing] 9.370 0.429 0.130 0.477 13.659 0.558 0.091 0.316
DeepSNNet [deng2025deepsn] 10.195 0.411 0.113 0.534 17.394 0.576 0.075 0.303
AWRaCLe [rajagopalan2025awracle] 11.208 0.431 0.091 0.513 10.540 0.495 0.121 0.331
MaIR [li2025mair] 11.072 0.423 0.086 0.529 23.637 0.770 0.005 0.271
@@
Figure 7: Visualization results of image translation (top row) and image inpainting (bottom row). Zoom in for best view.
@@ 5.5 Zero-Short Real-world Generation
Unknown task generalization. POLED and TOLED [zhou2021image] are captured by under-display cameras in high-resolution with different degradation types, which fully meet the real-world scene. The quantitative results are reported in Tab. 7 while the visual comparisons are illustrated in Fig. 6. Evidently, our method achieves the best metric evaluation and our restored image is the most similar to ground-truth.
Known task generalization. As real-world datasets mainly have no ground truth, we use the non-reference metric, i.e., MetaIQA [zhu2020metaiqa] and NIQE [mittal2012making], to assess the perceptual quality, as provided in Tab. 8. Results show that our method outperforms other universal models in various benchmarks.
@@
Restomer [zamir2022restormer] 0.231 13.115 0.328 3.828 0.357 5.845 0.437 4.400 0.303 6.734
AirNet [li2022all] 0.232 11.668 0.280 3.674 0.347 6.091 0.440 4.623 0.286 6.393
Prompt-IR [potlapalli2023promptir] 0.232 11.439 0.308 3.797 0.361 5.840 0.437 4.962 0.286 6.670
ProRes [ma2023prores] 0.226 13.110 0.348 3.933 0.355 5.976 0.434 5.444 0.297 6.574
IDR [zhang2023ingredient] 0.231 11.100 0.324 3.866 0.363 5.850 0.453 4.634 0.300 6.683
IRSDE [luo2023image] 0.230 11.391 0.351 3.809 0.357 5.874 0.427 4.134 0.285 6.289
AutoDIR [jiang2024autodir] 0.231 10.800 0.366 3.910 0.373 5.831 0.470 9.881 0.308 6.493
DA-CLIP [luo2024controlling] 0.232 10.604 0.334 3.720 0.361 5.864 0.460 6.531 0.310 6.058
GOUB [yue2024image] 0.231 11.566 0.373 3.928 0.360 5.853 0.458 4.104 0.278 6.303
ConvIR [cui2024revitalizing] 0.236 10.280 0.370 3.723 0.364 5.813 0.446 4.645 0.313 6.465
DeepSNNet [deng2025deepsn] 0.231 11.446 0.348 3.896 0.367 5.882 0.436 4.662 0.301 6.525
AWRaCLe [rajagopalan2025awracle] 0.232 12.016 0.366 3.796 0.363 5.898 0.426 4.649 0.306 6.516
MaIR [li2025mair] 0.234 10.804 0.350 3.666 0.363 5.890 0.245 22.446 0.284 6.590
@@ 5.7 Image Translation and Inpainting
RDBM owns distinct advantages in mapping the data distribution to the prior distribution, thereby enabling extensive validation on similar computer vision tasks. To this end, we expand our experimental settings on image-to-image translation and image inpainting to fully demonstrate the potential of our method. The former aims to transform an input image from one domain to another while preserving certain essential semantic or structural features. The latter focuses on filling in missing or corrupted regions within an image. Specifically, we adopt the widely used edge2handlebags [isola2017image] dataset for image-to-image translation and apply the Celebrate-HQ dataset [karras2018progressive] with masks provided in  [liu2018image] for image inpainting. All these datasets are scaled to 256 × 256. We additionally employ Fréchet Inception Distance (FID) [heusel2017gans] for evaluation. Qualitative comparisons and quantitative results are presented in Fig. 7 and Tab. 9, respectively. Clearly, our method achieves the best visual effects and the best metrics assessments.
@@
Method Image Translation [isola2017image] Image Inpainting [karras2018progressive]
@@
DDPM [songscore] 8.39 0.447 8.39 0.412 19.22 0.746 0.526 0.126
ReFlow [liuflow] 10.46 0.442 5.76 0.374 23.53 0.822 0.307 0.149
BBDM [li2023bbdm] 14.75 0.635 7.46 0.248 20.36 0.653 0.386 0.169
I2SB [liu20232] 12.57 0.615 6.64 0.357 27.34 0.890 0.379 0.054
RDDM [liu2024residual] 14.66 0.645 5.72 0.256 23.94 0.852 0.167 0.119
GOUB [yue2024image] 16.58 0.700 8.76 0.288 31.56 0.920 0.321 0.065
@@ Appendix A Doob’s $h$ transform
Proof: In theory, $p(\mathbf{x}_{t}\mid\mathbf{x}_{0})$ and $p(\mathbf{x}_{T}\mid\mathbf{x}_{t})$ satisfy the Kolmogorov Forward Equation (KFE) and Kolmogorov Backward Equation (KBE), respectively [risken1989fokker], as formulated below:
@@ Appendix E Process of Reverse Inference
$\mathbf{x}_{t-1}=\kappa_{t}\mathbf{x}_{t}+\eta_{t}\boldsymbol{\mu}+\gamma_{t}\mathbf{x}_{0}+\dot{\sigma_{t}}\boldsymbol{\pi}\epsilon_{t}$ (E.13)
$=\kappa_{t}(\boldsymbol{\mu}+(\mathbf{x}_{0}-\boldsymbol{\mu})\Theta_{t}+\boldsymbol{\pi}\Sigma_{t}\epsilon_{t})+\eta_{t}\boldsymbol{\mu}+\gamma_{t}\mathbf{x}_{0}+\dot{\sigma_{t}}\boldsymbol{\pi}\epsilon_{t}$
$=(\kappa_{t}+\eta_{t}-\kappa_{t}\Theta_{t})\boldsymbol{\mu}+(\kappa_{t}\Theta_{t}+\gamma_{t})\mathbf{x}_{0}+\boldsymbol{\pi}(\kappa_{t}^{2}\Sigma_{t}^{2}+\dot{\sigma_{t}}^{2})^{\frac{1}{2}}\epsilon_{t},$
@@
$\Sigma_{t-1}^{2}$ $=\kappa_{t}^{2}\Sigma_{t}^{2}+\dot{\sigma_{t}}^{2}.$ (E.16)
By setting $\dot{\sigma_{t}}=0$:
@@ Appendix F Connections Among Existing Diffusion Bridge
$d\mathbf{x}_{t}=\theta_{t}\coth(\overline{\theta}_{t:T})(\boldsymbol{\mu}-\mathbf{x}_{t})dt+\sqrt{2\boldsymbol{\pi}^{2}\lambda\theta_{t}}d\omega_{t},$ (F.2)
@@ F.1 Connections to Variance-Exploding (VE) and Variance-Preserving (VP) SDEs
SMLD [songscore] primarily introduces two mainstream diffusion formulations, namely VP and VE. For a given generalized OU process in Eq. (F.1), there exists relationships:
@@
where we set the $\theta_{t}\rightarrow\sigma_{t}^{2}$, implying $\lambda=\frac{1}{2}$. On this basis, DDBM [zhoudenoising] further extends such diffusion configuration to bridge models, which are also special cases of our formulation under specific configurations:
@@ F.2 Connections to Brownian Bridge SDEs
Brownian bridge is a fundamental architecture for diffusion model, which are widely adopted in BBDM [li2023bbdm], I2SB [liu20232]. By setting $\theta_{t}\rightarrow 0$ with condition $2\lambda\theta_{t}=1=\sigma_{t}^{2}$, we can derive the Brownian Bridge as formulated below:
@@
$=\lim\limits_{\theta_{t}\rightarrow 0,\sigma_{t}^{2}\rightarrow 1}^{\boldsymbol{\pi}=1}\{d\mathbf{x}_{t}=\frac{\boldsymbol{\mu}-\mathbf{x}_{t}}{T-t}dt+d\omega_{t}\}$
@@ F.3 Connections to Flow Matching
Flow-based generative models [lipman2023flow, liuflow] design a deterministic probability path that linearly interpolates between a prior and the data distribution, and then directly learn a time-dependent vector field whose integral trajectories realize this path. By discarding the stochastic noise ($\sigma_{t}=0$) and adopting the Brownian bridge configuration, Eq. (F.2) can be transformed into:
@@ F.4 Connections to OU Bridge SDEs
Eq. (F.2) can be transformed into naive OU bridge [yue2024image] by setting $\pi=1$ to recover global noise perturbation:
@@ F.5 Connections to Stochastic Interpolants
Stochastic interpolants [albergo2023stochastic] define a unified framework for flows and diffusions, which can be expressed as:
@@ H.1 Summary about the Datasets
Table S1: Summary of the image restoration datasets utilized in this paper.
@@
Deraining DID [zhang2018density] Synthetic - 1,200
Rain13K [Kui_2020_CVPR] Synthetic 13,711 -
Rain_100 [yang2017deep] Synthetic - 200
DeRaindrop [qian2018attentive] Real 861 307
GT-Rain [ba2022gt-rain] Real 26,125 2,100
RealRain-1k [li2022toward] Real 1792 448
Low-light Enhancement LOL [wei2018deep] Real 485 15
@@
VE-LOL-L [ll_benchmark] Synthetic/Real 900/400 100/100
NPE [wang2013naturalness] Real - 8
DICM [lee2013contrast] Real - 64
Desnowing CSD [chen2021all] Synthetic 8,000 2,000
Snow100K-Real [liu2018desnownet] Real - 1,329
Dehazing SOTS [li2018benchmarking] Synthetic - 500
ITS_v2 [li2018benchmarking] Synthetic 13,990 -
D-HAZY [Ancuti_D-Hazy_ICIP2016] Synthetic 1,178 294
NH-HAZE [NH-Haze_2020] Real - 55
Dense-Haze [ancuti2019dense] Real - 55
NHRW [zhang2017fast] Real - 150
Deblur GoPro [nah2017deep] Synthetic 2,103 1,111
RealBlur [rim_2020_ECCV] Real 3,758 980
@@ H.3 More Visual Comparisons on Real-world Scene Generalization
Unknown task generalization. Unknown task image restoration is performed on both POLED and TOLED [zhou2021image]. Visual comparisons on the POLED dataset are provided in Fig. S10. The results show that our method can generalize to real-world scenes and achieve competitive visual results.
Figure S10: Visualization results of zero-shot generalization in real-world POLED dataset. Zoom in for best view.
@@ H.4 More Visual Comparisons on Image Translation and Inpainting
Figure S11: Visualization results of image translation. Zoom in for best view.
Figure S12: Visualization results of image inpainting. Zoom in for best view.
H.5 Efficiency Comparison
Our mixed dataset consists of images with resolutions ranging from 256 to 1024 pixels. Accordingly, we evaluate model efficiency under three representative resolution settings, as summarized in Tab. S2. Evidently, our methods are moderately efficient with reasonable resource consumption. Overall, RDBM strikes a balance between efficiency and performance.
Table S2: Efficiency comparisons among universal methods. ’-’ means out of memmory.
Resolution $\times256$ $\times512$ $\times1024$
Method Mem.(G) Time(s) FPS Mem.(G) Time(s) FPS Mem.(G) Time(s) FPS
Restomer [zamir2022restormer] 1.959 0.105 9.563 6.670 0.381 2.622 25.419 1.773 0.564
AirNet [li2022all] 1.039 0.194 5.159 3.480 0.738 1.355 11.244 20.499 0.049
Prompt-IR [potlapalli2023promptir] 2.544 0.111 8.981 7.255 0.399 2.508 26.005 1.845 0.542
ProRes [ma2023prores] 2.027 0.318 3.149 2.514 0.766 1.305 6.025 1.715 0.583
IDR [zhang2023ingredient] 1.340 0.052 19.253 4.313 0.136 7.373 16.110 0.615 1.626
IRSDE [luo2023image] 1.554 5.017 0.199 2.743 18.493 0.054 9.997 72.289 0.014
AutoDIR [jiang2024autodir] 7.023 6.266 0.160 11.021 11.986 0.083
DA-CLIP [luo2024controlling] 2.119 2.585 0.387 6.775 7.937 0.126 58.548 60.893 0.016
GOUB [yue2024image] 1.554 4.996 0.200 2.868 18.442 0.054 10.122 72.239 0.014
ConvIR [cui2024revitalizing] 0.708 0.035 28.570 1.184 0.055 18.020 2.809 0.192 5.202
DeepSNNet [deng2025deepsn] 0.862 0.100 9.974 0.989 0.102 9.801 1.364 0.267 3.749
AWRaCLe [rajagopalan2025awracle] 1.929 0.101 9.922 4.264 0.354 2.822 13.608 1.374 0.728
MaIR [li2025mair] 2.091 1.297 0.771 6.593 4.744 0.211 24.593 18.029 0.055
RDBM-T 0.813 0.418 2.392 1.907 1.621 0.617 13.407 6.648 0.150
RDBM-S 0.825 0.431 2.322 1.921 1.648 0.607 13.421 6.775 0.148
RDBM-B 1.124 0.480 2.081 2.186 1.926 0.519 14.938 7.872 0.127
RDBM-L 1.150 0.504 1.986 2.307 1.982 0.505 15.059 8.135 0.123